# Do Language Models Associate Sound with Meaning? A Multimodal Study of Sound Symbolism

### v4

@@ 1 Introduction
Sound symbolism, which suggests that phonetic sounds and their meanings have a significant correlation, presents a cognitively grounded exception to the principle of linguistic arbitrariness (hinton2006sound; dingemanse2015arbitrariness). For instance, when people are presented with two images of sharp and round shapes called ``kiki'' and ``bouba'', the overwhelming majority of participants consistently match ``kiki'' with pointed shapes and ``bouba'' with round shapes, regardless of their cultural background (kohler1967gestalt; ramachandran2001synaesthesia). This iconicity effect illustrates the universal sensory associations of humans that intuitively link certain phonetic sounds with certain semantic features, which facilitate early-childhood or second-language acquisition (imai2014sound), and commercially nuanced brand naming (yorkston2004sound).
Recent advancements in Multimodal Large Language Models (MLLMs) that incorporate audio-modality input into an integrated representation space (openai2024gpt4ocard; xu2025qwen25omnitechnicalreport) shed light on new possibilities to systematically analyze human-like phonetic iconicity of language models. With MLLMs as a novel test bed, we formulate the following two key research questions:
@@
Mimetic words provide a compelling probe to address these subjects. Represented by onomatopoeias and ideophones [^1] In this paper, “onomatopoeia” and “ideophone” refer to words that depict sounds and non-auditory sensory imagery, respectively., mimetic words refer to words in which the non-arbitrary association of sound form and meaning is prominent (akita2019ideophones), such as ``boom'' or ``whizz.'' We construct LEX-ICON, a dataset of natural mimetic words and constructed pseudo-words designed to maximize sound-symbolic effect. With LEX-ICON, we apply the semantic dimension paradigm, where word meanings are projected onto binary scales (e.g., fast vs. slow), to examine MLLMs' abilities to infer the meaning of words from their form (original text, IPA, audio) rather than their content. We further analyze models' internal representations to identify whether models actually focus on iconic phonemes, as illustrated in Figure 1.
@@ 2.1 Sound Symbolism in Linguistics
Early modern linguists question the arbitrary relationship between signifier and signified, suggesting the phenomenon of intuitive phonetic symbolism in humans (usnadze1924experimenteller; sapir1929study; kohler1967gestalt). sapir1929study's experiment has exhibited that people who are presented with unfamiliar object names ``mil'' and ``mal'' and are asked to estimate the size of the objects respond that the latter is larger than the former (parise2012audiovisual). Recent studies continue to observe these phenomena (hinton2006sound; lockwood2015iconicity; cwiek2022bouba), extending them by scaling to a large amount of data (thompson2021articulatory; winter2024iconicity) or measuring the exact semantic dimensions associated with each phoneme (monaghan2019sound; sidhu2022higher; sidhu2025sound). For the mimetic words, experimental research has also demonstrated that people can infer a certain degree of meaning from the words in languages besides their mother tongue (shinohara2010cross; dingemanse2016sound).
@@ 2.2 Phonetic Iconicity for LLMs
Iconicity tasks for language models have been used as a means of assessing whether the models have human-like phonetic intuition (cai-etal-2024-large; duan2024hlbbenchmarkingllmshumanlikeness). Early studies focus on analyzing phonesthemic information contained in word embedding spaces (abramova2013automatic; abramova-fernandez-2016-questioning). Recent studies demonstrate that the effect of non-arbitrary integration of linguistic form and meaning is substantiated in text-only LLMs (miyakawa-etal-2024-llms; marklova2025iconicitylargelanguagemodels), vision and image generation models (loakman-etal-2024-ears; alper2024kikiboubasoundsymbolism; shinto2024analyzing; iida2024investigating), and audio-visual models (tseng2024measuring), yet they are not extensive enough in scope to include exhaustive and multilingual mimetic word data or diverse semantic feature analysis.
@@ 2.3 Multimodal Interpretability
Recently, the field of model interpretability has witnessed significant growth (elhage2021mathematical; zou2025representationengineeringtopdownapproach; wang2023interpretability). Motivated by the rapid progress of MLLMs (openai2024gpt4ocard; xu2025qwen25omnitechnicalreport), there has been increasing interest in multimodal interpretability (lin2025surveymechanisticinterpretabilitymultimodal). For instance, neo2025towards ablates a subset of visual tokens to observe the resulting differences in model output, while nikankin2025taskdifferentcircuitsdisentangling demonstrates that models employ distinct processing circuits depending on the input modality. Despite these advances, prior work has predominantly focused on the visual modality, with less exploration of audio-based interpretability. yang2025audiolenscloserlookauditory investigates how models handle auditory input, but their analysis is limited to simple settings where the audio simply serves as a direct vocalization of textual data. In this work, we examine audio interpretability through the lens of sound symbolism, highlighting how phoneme-level features encode meaning in ways that are intrinsically tied to the auditory modality.
@@ 3 LEX-ICON
Table 1: Data distribution of LEX-ICON across natural (8,052 words from English, French, Japanese, and Korean) and constructed (2,930 words) groups. Japanese and Korean are known for their rich mimetic words (hamano1986sound; kwon2018iconicity), which contributes to their high proportion.
@@
Table 2: Semantic dimension distribution of pseudo ground truth data by word group. We adopt 25 semantic dimension criteria by sidhu2022higher to annotate LEX-ICON. Six dimensions from the constructed word group are excluded by removing close-to-neutral data points. As a result, 19 dimensions remain for experiments in §4.
@@ 3.1 Semantic Dimension
We employ the semantic dimension methodology, a concept originated from ``semantic differential'' by osgood1957measurement, which consists of two semantic feature adjectives located at opposite extremes, such as ``big'' and ``small.'' By applying this methodology to sound symbolism experiments, we can simultaneously and precisely measure the multifactorial meanings contained in a single word. We annotate up to 25 pairs of predefined semantic features as per sidhu2022higher for each word in the LEX-ICON, which is the most diverse scale in terms of studies on LLMs' iconicity.
@@ 3.2 Natural Mimetic Words
We manually collect 8,052 mimetic words and definition data consisting of onomatopoeias and ideophones in English, French, Japanese, and Korean from specialized mimetic word lexicons and authoritative dictionaries for each language. We then extract the most representative definitions of these mimetic words from dictionaries such as the Oxford English Dictionary (dictionary1989oxford), Le Petit Robert (lepetitrobert2023), Nihon Kokugo Daijiten (9784095210216), and the Standard Korean Language Dictionary (nkl2025). For further source information, see the Appendix.
@@
To observe the effect of trained token memorization and modality change, we create three types of word form that contain textual and auditory input: (1) original text, (2) IPA-converted text with phoneme-level spacing, (3) audio waveform converted using text-to-speech (TTS) software. We perform IPA conversion with the Epitran package (Mortensen-et-al:2018), and obtain the TTS dataset using Google Text-to-Speech (google_tts_api_2025) for English, French, and Japanese; and MeloTTS (zhao2024melo) for Korean.
@@
To effectively generate a large-scale ground truth data, all natural language words and their definitions in the LEX-ICON are given four LLMs: GPT-4.1 (openai2025_gpt41), Qwen3-32B (yang2025qwen3technicalreport) [^2] Non-reasoning mode., Gemma-3-27B (gemmateam2025gemma3technicalreport), and Gemini-2.5-flash (comanici2025gemini25pushingfrontier). The models are prompted to annotate each semantic dimension of a word with one of a feature pair or neutral labels (e.g., selecting one of the ``exciting'', ``calming'', or ``neither'' option) for each word. See the Appendix for the detailed prompt.
@@ 3.3 Constructed Mimetic Words
We systematically construct novel words using a CVCV structure to create pseudo-words unlikely to be encountered during model training. We use 15 consonants from five categories: sonorants (/l/, /m/, /n/), voiced fricatives (/v/, /D/, /z/), voiceless fricatives (/f/, /s/, /S/), voiceless stops (/p/, /t/, /k/), and voiced stops (/b/, /d/, /g/). We use four vowels from two categories: front vowels (/i/, /ej/) and back vowels (/A/, /ow/). After filtering against existing entries in the IPA-dict database (ipa_dict) across four languages (US and UK English, French, Japanese, and Korean), we obtain 3,108 novel pseudo-words. We convert IPA symbols to English alphabet combinations for TTS compatibility and remove incorrectly pronounced words by Google TTS, yielding 2,930 final words.
@@
We utilize the empirical coefficients from sidhu2022higher, who quantified associations between phoneme categories and semantic dimensions through human rating experiments. These coefficients represent how much phoneme categories deviate from the overall mean rating across 25 semantic dimensions. We assign corresponding scores to phonemes in our constructed words and calculate the mean score across the four phonemes. To focus on meaningful associations, we apply a threshold of 1.0 standard deviation from the neutral point, treating data points within this range and as semantically neutral, consistent with 'neither' labels used for multilingual mimetic words.
@@ 4.1 Experimental Settings
We employ MLLMs that officially support simultaneous text and audio inputs: Qwen2.5-Omni (xu2025qwen25omnitechnicalreport) [^3] Qwen2.5-Omni-3B and Qwen2.5-Omni-7B., Gemini-2.5-flash (geminiteam2025geminifamilyhighlycapable), and GPT-4o (openai2024gpt4ocard)444gpt-4o for text-only input, gpt-4o-audio-preview for audio-enabled input.. All experimental models cover the four languages and orthographies present in LEX-ICON. For the Qwen models, inference is performed on one RTX 4090 GPU. We apply a zero-shot prompting strategy with temperature set to 0 throughout all experiments to ensure reproducibility.
@@ 4.3 Result
In Figure 3, MLLMs' macro-F1 scores averaged across all input types surpass the baseline score (0.50) in 84.2% (natural group) and 68.4% (constructed group) of the semantic dimensions, indicating that the models can detect phonetic iconicity not only in natural mimetic words that may have been memorized during training phase, but also in constructed pseudo-words with maximized sound-symbolic effects. Overall performance becomes even larger when the comparatively small-scaled Qwen2.5-Omni-3B model is excluded. These results are supported by human evaluation results that score above the baseline in most dimensions, which guarantees the reliability of our pseudo ground truth data automatically annotated by LLMs from dictionary data in §3.1. Notably, the models' strong performance on the sharp vs. round dimension aligns with well-known cognitive linguistic experiments like the ``bouba-kiki'' effect (ramachandran2001synaesthesia).
@@
Our findings reveal systematic modality preferences across semantic dimensions, providing computational evidence for the multi-mechanism nature of sound symbolism, as shown in Figure 5. For dimensions where acoustic features are theoretically central, such as size distinctions (big vs. small) that correlate with formant frequencies (knoeferle2017drives) and speed distinctions (fast vs. slow) that relate to consonant voicing duration (saji2013cross), the MLLMs show a pattern of enhanced performance when processing constructed words in audio format, consistent with the theoretical predictions. Conversely, for dimensions where non-acoustic mechanisms are proposed, such as shape associations (sharp vs. round) based on lip rounding gestures (imai2025does) and valence associations (beautiful vs. ugly, happy vs. sad) affected by articulatory properties (korner2022articulation), the models exhibit stronger reliance on textural representations.
@@ 5 Internal Attention Analysis
Figure 6: Attention fraction scores indicating the ratio of semantic dimensions that the model focuses on for each IPA, by word group and input type. For each IPA and semantic dimension, the model tends to attend more to semantic features that exhibit stronger associations with the given phonemes in linguistics. For instance, the model mostly associates sharp semantic feature with /p/ and /k/, round with /m/ and /n/, big with /A/, and small with /i/ (kohler1967gestalt; parise2012audiovisual).
@@ 5.2 Methodology
When an IPA symbol spans multiple text tokens, the procedure first sums the attention scores across those tokens before fraction normalization. For audio inputs, we employ Montreal Forced Aligner (mcauliffe2025montreal) to segment each audio waveform into phonemes over time. The phoneme sequence is then aligned with the model’s 40 ms sampling period, with consecutive occurrences of the same phoneme treated as a single IPA symbol. Further details are in the Appendix.
@@ 5.3 Result
Figure 6 presents heatmaps of attention fraction scores for canonical IPA symbols and semantic dimensions by input type and word group. The patterns for constructed words in the IPA modality closely mirror prior findings in sound symbolism research. For example, phonemes such as /p/ and /k/ exhibit elevated attention under the sharp feature, whereas /m/ and /n/ associate with the round feature (kohler1967gestalt).
@@ A.1 Mimetic Word Data
We manually extract headwords and definitions from published onomatopoeia dictionaries for each language to determine whether a word belongs to the mimetic word category. For Korean, the word definition guideline of the NIKL_StandardKoreanLanguageDictionaryCompilationGuideline2_2000 is used to extract mimetic words and definitions. Semantic definitions that need to be supplemented are obtained from authoritative dictionaries for each language. The full list of the dictionaries is provided in Table 4.
@@
English Rhyme over Reason: Phonological Motivation in English (benczes2019rhyme)
Oxford English Dictionary, 2nd Edition (dictionary1989oxford)
French Dictionnaire des Onomatopées (enckell2003dictionnaire)
Le Petit Robert de la Langue Française (lepetitrobert2023)
Trésor de la Langue Française (letresor1994)
Japanese Kurashi no Kotoba Gion Gitaigo Jiten (1130282271732203008)
Seisenban Nihon Kokugo Daijiten (9784095210216)
Kojien, 7th Edition (kojien7)
Digital Daijisen (digitaldaijisen)
Daijirin, 4th Edition (matsumura2019daijirin)
Korean Standard Korean Language Dictionary (nkl2025)

### v5

@@ 1 Introduction
Sound symbolism, which suggests that phonetic sounds and their meanings have a significant correlation, presents a cognitively grounded exception to the principle of linguistic arbitrariness (Hinton and others 2006; Dingemanse et al. 2015). For instance, when people are presented with two images of sharp and round shapes called ``kiki'' and ``bouba'', the overwhelming majority of participants consistently match ``kiki'' with pointed shapes and ``bouba'' with round shapes, regardless of their cultural background (Köhler 1967; Ramachandran and Hubbard 2001). This iconicity effect illustrates the universal sensory associations of humans that intuitively link certain phonetic sounds with certain semantic features, which facilitate early-childhood or second-language acquisition (Imai and Kita 2014), and commercially nuanced brand naming (Yorkston and Menon 2004).
Recent advancements in Multimodal Large Language Models (MLLMs) that incorporate audio-modality input into an integrated representation space (OpenAI et al. 2024; Xu et al. 2025) shed light on new possibilities to systematically analyze human-like phonetic iconicity of language models. With MLLMs as a novel test bed, we formulate the following two key research questions:
@@
Mimetic words provide a compelling probe to address these subjects. Represented by onomatopoeias and ideophones [^1] In this paper, “onomatopoeia” and “ideophone” refer to words that depict sounds and non-auditory sensory imagery, respectively., mimetic words refer to words in which the non-arbitrary association of sound form and meaning is prominent (Akita and Pardeshi 2019), such as ``boom'' or ``whizz.'' We construct LEX-ICON, a dataset of natural mimetic words and constructed pseudo-words designed to maximize sound-symbolic effect. With LEX-ICON, we apply the semantic dimension paradigm, where word meanings are projected onto binary scales (e.g., fast vs. slow), to examine MLLMs' abilities to infer the meaning of words from their form (original text, IPA, audio) rather than their content. We further analyze models' internal representations to identify whether models actually focus on iconic phonemes, as illustrated in Figure 1.
@@ 2.1 Sound Symbolism in Linguistics
Early modern linguists question the arbitrary relationship between signifier and signified, suggesting the phenomenon of intuitive phonetic symbolism in humans (Usnadze 1924; Sapir 1929; Köhler 1967). Sapir (1929)'s experiment has exhibited that people who are presented with unfamiliar object names ``mil'' and ``mal'' and are asked to estimate the size of the objects respond that the latter is larger than the former (Parise and Spence 2012). Recent studies continue to observe these phenomena (Hinton and others 2006; Lockwood and Dingemanse 2015; Ćwiek et al. 2022), extending them by scaling to a large amount of data (Thompson et al. 2021; Winter et al. 2024) or measuring the exact semantic dimensions associated with each phoneme (Monaghan and Fletcher 2019; Sidhu et al. 2022; Sidhu 2025). For the mimetic words, experimental research has also demonstrated that people can infer a certain degree of meaning from the words in languages besides their mother tongue (Shinohara and Kawahara 2010; Dingemanse et al. 2016).
@@ 2.2 Phonetic Iconicity for LLMs
Iconicity tasks for language models have been used as a means of assessing whether the models have human-like phonetic intuition (Cai et al. 2024; Duan et al. 2024). Early studies focus on analyzing phonesthemic information contained in word embedding spaces (Abramova et al. 2013; Abramova and Fernández 2016). Recent studies demonstrate that the effect of non-arbitrary integration of linguistic form and meaning is substantiated in text-only LLMs (Miyakawa et al. 2024; Marklová et al. 2025), vision and image generation models (Loakman et al. 2024; Alper and Averbuch-Elor 2024; Shinto and Iizuka 2024; Iida and Funakura 2024), and audio-visual models (Tseng et al. 2024), yet they are not extensive enough in scope to include exhaustive and multilingual mimetic word data or diverse semantic feature analysis.
@@ 2.3 Multimodal Interpretability
Recently, the field of model interpretability has witnessed significant growth (Elhage et al. 2021; Zou et al. 2025; Wang et al. 2023). Motivated by the rapid progress of MLLMs (OpenAI et al. 2024; Xu et al. 2025), there has been increasing interest in multimodal interpretability (Lin et al. 2025). For instance, Neo et al. (2025) ablates a subset of visual tokens to observe the resulting differences in model output, while Nikankin et al. (2025) demonstrates that models employ distinct processing circuits depending on the input modality. Despite these advances, prior work has predominantly focused on the visual modality, with less exploration of audio-based interpretability. Yang et al. (2025b) investigates how models handle auditory input, but their analysis is limited to simple settings where the audio simply serves as a direct vocalization of textual data. In this work, we examine audio interpretability through the lens of sound symbolism, highlighting how phoneme-level features encode meaning in ways that are intrinsically tied to the auditory modality.
@@ 3 LEX-ICON
Table 1: Data distribution of LEX-ICON across natural (8,052 words from English, French, Japanese, and Korean) and constructed (2,930 words) groups. Japanese and Korean are known for their rich mimetic words (Hamano 1986; Kwon 2018), which contributes to their high proportion.
@@
Table 2: Semantic dimension distribution of pseudo ground truth data by word group. We adopt 25 semantic dimension criteria by Sidhu et al. (2022) to annotate LEX-ICON. Six dimensions from the constructed word group are excluded by removing close-to-neutral data points. As a result, 19 dimensions remain for experiments in §4.
@@ 3.1 Semantic Dimension
We employ the semantic dimension methodology, a concept originated from ``semantic differential'' by Osgood et al. (1957), which consists of two semantic feature adjectives located at opposite extremes, such as ``big'' and ``small.'' By applying this methodology to sound symbolism experiments, we can simultaneously and precisely measure the multifactorial meanings contained in a single word. We annotate up to 25 pairs of predefined semantic features as per Sidhu et al. (2022) for each word in the LEX-ICON, which is the most diverse scale in terms of studies on LLMs' iconicity.
@@ 3.2 Natural Mimetic Words
We manually collect 8,052 mimetic words and definition data consisting of onomatopoeias and ideophones in English, French, Japanese, and Korean from specialized mimetic word lexicons and authoritative dictionaries for each language. We then extract the most representative definitions of these mimetic words from dictionaries such as the Oxford English Dictionary (Simpson 1989), Le Petit Robert (Alain Rey 2022), Nihon Kokugo Daijiten (Shogakukan 2006), and the Standard Korean Language Dictionary (National Institute of Korean Language 2025). For further source information, see the Appendix.
@@
To observe the effect of trained token memorization and modality change, we create three types of word form that contain textual and auditory input: (1) original text, (2) IPA-converted text with phoneme-level spacing, (3) audio waveform converted using text-to-speech (TTS) software. We perform IPA conversion with the Epitran package (Mortensen et al. 7–12), and obtain the TTS dataset using Google Text-to-Speech (Google 2025) for English, French, and Japanese; and MeloTTS (Zhao et al. 2023) for Korean.
@@
To effectively generate a large-scale ground truth data, all natural language words and their definitions in the LEX-ICON are given four LLMs: GPT-4.1 (OpenAI 2025), Qwen3-32B (Yang et al. 2025a) [^2] Non-reasoning mode., Gemma-3-27B (Team et al. 2025b), and Gemini-2.5-flash (Comanici et al. 2025). The models are prompted to annotate each semantic dimension of a word with one of a feature pair or neutral labels (e.g., selecting one of the ``exciting'', ``calming'', or ``neither'' option) for each word. See the Appendix for the detailed prompt.
@@ 3.3 Constructed Mimetic Words
We systematically construct novel words using a CVCV structure to create pseudo-words unlikely to be encountered during model training. We use 15 consonants from five categories: sonorants (/l/, /m/, /n/), voiced fricatives (/v/, /D/, /z/), voiceless fricatives (/f/, /s/, /S/), voiceless stops (/p/, /t/, /k/), and voiced stops (/b/, /d/, /g/). We use four vowels from two categories: front vowels (/i/, /ej/) and back vowels (/A/, /ow/). After filtering against existing entries in the IPA-dict database (Doherty 2016) across four languages (US and UK English, French, Japanese, and Korean), we obtain 3,108 novel pseudo-words. We convert IPA symbols to English alphabet combinations for TTS compatibility and remove incorrectly pronounced words by Google TTS, yielding 2,930 final words.
@@
We utilize the empirical coefficients from Sidhu et al. (2022), who quantified associations between phoneme categories and semantic dimensions through human rating experiments. These coefficients represent how much phoneme categories deviate from the overall mean rating across 25 semantic dimensions. We assign corresponding scores to phonemes in our constructed words and calculate the mean score across the four phonemes. To focus on meaningful associations, we apply a threshold of 1.0 standard deviation from the neutral point, treating data points within this range and as semantically neutral, consistent with 'neither' labels used for multilingual mimetic words.
@@ 4.1 Experimental Settings
We employ MLLMs that officially support simultaneous text and audio inputs: Qwen2.5-Omni (Xu et al. 2025) [^3] Qwen2.5-Omni-3B and Qwen2.5-Omni-7B., Gemini-2.5-flash (Team et al. 2025a), and GPT-4o (OpenAI et al. 2024)444gpt-4o for text-only input, gpt-4o-audio-preview for audio-enabled input.. All experimental models cover the four languages and orthographies present in LEX-ICON. For the Qwen models, inference is performed on one RTX 4090 GPU. We apply a zero-shot prompting strategy with temperature set to 0 throughout all experiments to ensure reproducibility.
@@ 4.3 Result
In Figure 3, MLLMs' macro-F1 scores averaged across all input types surpass the baseline score (0.50) in 84.2% (natural group) and 68.4% (constructed group) of the semantic dimensions, indicating that the models can detect phonetic iconicity not only in natural mimetic words that may have been memorized during training phase, but also in constructed pseudo-words with maximized sound-symbolic effects. Overall performance becomes even larger when the comparatively small-scaled Qwen2.5-Omni-3B model is excluded. These results are supported by human evaluation results that score above the baseline in most dimensions, which guarantees the reliability of our pseudo ground truth data automatically annotated by LLMs from dictionary data in §3.1. Notably, the models' strong performance on the sharp vs. round dimension aligns with well-known cognitive linguistic experiments like the ``bouba-kiki'' effect (Ramachandran and Hubbard 2001).
@@
Our findings reveal systematic modality preferences across semantic dimensions, providing computational evidence for the multi-mechanism nature of sound symbolism, as shown in Figure 5. For dimensions where acoustic features are theoretically central, such as size distinctions (big vs. small) that correlate with formant frequencies (Knoeferle et al. 2017) and speed distinctions (fast vs. slow) that relate to consonant voicing duration (Saji et al. 2013), the MLLMs show a pattern of enhanced performance when processing constructed words in audio format, consistent with the theoretical predictions. Conversely, for dimensions where non-acoustic mechanisms are proposed, such as shape associations (sharp vs. round) based on lip rounding gestures (Imai et al. 2025) and valence associations (beautiful vs. ugly, happy vs. sad) affected by articulatory properties (Körner and Rummer 2022), the models exhibit stronger reliance on textural representations.
@@ 5 Internal Attention Analysis
Figure 6: Attention fraction scores indicating the ratio of semantic dimensions that the model focuses on for each IPA, by word group and input type. For each IPA and semantic dimension, the model tends to attend more to semantic features that exhibit stronger associations with the given phonemes in linguistics. For instance, the model mostly associates sharp semantic feature with /p/ and /k/, round with /m/ and /n/, big with /A/, and small with /i/ (Köhler 1967; Parise and Spence 2012).
@@ 5.2 Methodology
When an IPA symbol spans multiple text tokens, the procedure first sums the attention scores across those tokens before fraction normalization. For audio inputs, we employ Montreal Forced Aligner (McAuliffe and others 2025) to segment each audio waveform into phonemes over time. The phoneme sequence is then aligned with the model’s 40 ms sampling period, with consecutive occurrences of the same phoneme treated as a single IPA symbol. Further details are in the Appendix.
@@ 5.3 Result
Figure 6 presents heatmaps of attention fraction scores for canonical IPA symbols and semantic dimensions by input type and word group. The patterns for constructed words in the IPA modality closely mirror prior findings in sound symbolism research. For example, phonemes such as /p/ and /k/ exhibit elevated attention under the sharp feature, whereas /m/ and /n/ associate with the round feature (Köhler 1967).
@@ A.1 Mimetic Word Data
We manually extract headwords and definitions from published onomatopoeia dictionaries for each language to determine whether a word belongs to the mimetic word category. For Korean, the word definition guideline of the National Institute of Korean Language (2000) is used to extract mimetic words and definitions. Semantic definitions that need to be supplemented are obtained from authoritative dictionaries for each language. The full list of the dictionaries is provided in Table 4.
@@
English Rhyme over Reason: Phonological Motivation in English (Benczes 2019)
Oxford English Dictionary, 2nd Edition (Simpson 1989)
French Dictionnaire des Onomatopées (Enckell and Rézeau 2003)
Le Petit Robert de la Langue Française (Alain Rey 2022)
Trésor de la Langue Française (Centre national de la recherche scientifique and Institut national de la langue française 1994)
Japanese Kurashi no Kotoba Gion Gitaigo Jiten (Yamaguchi 2003)
Seisenban Nihon Kokugo Daijiten (Shogakukan 2006)
Kojien, 7th Edition (Shinmura 2018)
Digital Daijisen (Shogakukan 2025)
Daijirin, 4th Edition (Matsumura 2019)
Korean Standard Korean Language Dictionary (National Institute of Korean Language 2025)